# Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulations

### v2

@@ Acknowledgements
We thank the helpful discussions with Antonios, Katsianis and Cheng, Li. This work is supported by the National SKA Program of China (Grant Nos. 2022SKA0110200 and 2022SKA0110202), by the National Natural Science Foundation of China (NFSC) through grants 11733010 and 12173102, and by the China Manned Space Program through its Space Application System. The authors would like to thank National Supercomputer Center in Guangzhou for providing high performance computational resources. Post simulation analysis was performed on the HPC facility of the School of Physics and Astronomy, Sun Yat-Sen University.
@@ A.1 Stellar component
Observations in the 1970s and 1980s had identified M82 as a dusty disk galaxy with a central core. Later observations further revealed the existence of more complex structures. For instance, Telesco et al. (1991) identified characteristics of a bar in the central region of M82. The same structure was confirmed by Wills et al. (2000). In addition, some studies revealed the existence of spiral arms (Mayya et al. 2005) in M82. However, in this work, we omit the effect of bar and spiral arms, only account for a stellar fraction from a bulge, or in some contexts stellar spheroid (Tomisaka & Bregman 1993; Strickland & Stevens 2000), and an exponential disk. The scale height and scale length of both the bulge and stellar disk have been constrained by several observations (Ichikawa et al. 1995; Schreiber et al. 2001;Mayya et al. 2006).
@@
To describe the mass distribution and the potential of the stellar disk component, we use the widely adopted Miyamoto & Nagai (1975) disk profile. The density profile of a Miyamoto-Nagai disk is described by
@@
The bulge overlaps spatially with the recent starburst region. Therefore, a significant portion of the bulge’s stellar mass likely originated from the recent starburst events we aim to simulate. Hence, unlike previous studies, we take a different approach in our simulation by incorporating the bulge as a part of the initial gas disk. By directly adding the mass of the bulge to the initial mass of the gas disk, we can obtain a gas disk profile that more accurately reflect the initial condition before the onset of the recent starburst. This method also allows for a more self-consistent formation of stars in the nuclear region during our starburst simulation. The detailed setup of the gas disk is described in the following section.
Note that, there are some uncertainties associated with the chosen parameters for both the bulge and stellar disk. These uncertainties arise from the notable extinction and assumptions on the stellar population, initial mass function, star formation history, etc. A more accurate understanding of the stellar component may be of considerable interest, but is beyond our ability.
@@ A.2 Gas disk
Most of the previous simulational work of M82 uses a single Miyamoto & Nagai (1975) gas disk (e.g., Strickland & Stevens 2000; Cooper et al. 2008). More recently, Martini et al. (2018) proposed a different setup of the mass model to obtain the analytical form of the gravity potential. In their work, a triple Miyamoto-Nagai disk model (Smith et al. 2015) was employed. Though the Smith et al. (2015) model may produce an accurate fit to the mass distribution of an exponential gas disk, it fails to reproduce the gas profile of M82. Leroy et al. (2015) found the density profile of HI in M82 best fitted by a projected $\rho\propto r^{-2}$ density profile, and $\rm{H_{2}}$ best fitted by a projected $\rho\propto r^{-3.5}$ density profile along the major axis. We employ a similar method to that described by Smith et al. (2015), using three Miyamoto-Nagai disks to fit the projected density profile found by Leroy et al. (2015). Our choice of parameters for the triple Miyamoto-Nagai disk model is given in Table 3.
@@ A.3 Dark matter halo
Early research of the M82’s rotational curve indicated that its rotational curve profile was likely to be Keplerian (Goetz et al. (1990)). Consequently, many previous simulations of M82 did not include the gravitational contribution from the dark matter halo. However, more recent observations of the CO line and star clusters suggest a different scenario. Greco et al. (2012) found the rotational curve of M82 is considerably flattened towards the outer region, which is commonly interpreted as evidence for the existence of the dark matter halo (DM halo hereafter). However, the complexity of interactions and mergers among the three members of the M81 group complicates efforts to accurately constrain the parameters of the DM halo. In Oehm et al. (2017), three-body simulation and Markov chain Monte Carlo (MCMC) method are employed to constrain the initial parameters of the three DM halos, and high-resolution N-body simulation was conducted to further investigate the evolution of the triplet. They demonstrated that the current DM halos, in which M82 and its neighboring galaxies are embedded, arise from distortion, blending, and merging of the original DM halos. Simple analytical models centered at M82 may struggle to reproduce the true profile of the galaxy’s DM halo. For the sake of simplicity, we adopt a modified NFW profile in this work. The density distribution of the original NFW dark matter halo is given by
@@
Though the NFW profile has been widely adopted, it is not without its shortcomings, particularly in the scenario we are considering. The density of an NFW DM halo does not converge to a finite value towards the center, while observations suggest that many galaxies have a constant-density core in the center area of their DM halos. Such discrepancy is known as the cusp-core problem. Read et al. (2016) in their simulation shows that star formation and stellar feedback activities occurring in the embedded dwarf galaxy can transform the initial NFW DM halo into a ’core-NFW’ profile with a constant density core in the starburst region ($<R_{1/2}$) within a Hubble time. The profile of a ’Core-NFW’ halo can be described by
@@
With the $c-\rm{M_{vir}}$ relation enforced, two free parameters remain: $\rm{M_{vir}}$ and $R_{vir}$. We used the Markov Chain Monte Carlo(MCMC) method to find the optimal combination of the two parameters that best fit the observed rotation curve of Greco et al. (2012). To obtain a more accurate result, we excluded the central 200 pc region, where the velocity dispersion is significantly higher than the local rotational velocity. Additionally, we also excluded the velocity bump located $\sim 500$ pc at the eastern part of M82, which is caused by an SSC designated as cluster ’z’ by McCrady & Graham (2007). We find the best fitting halo mass to be $\rm{M_{vir}}=6.0\times 10^{10}\,\rm{M_{\odot}}$, and the best fitting virial radius is $R_{vir}=53\,\rm{kpc}$. Fig 1 shows that the overall rotation curve of our model closely matches the observed result of Greco et al. (2012).
Appendix B Convergence test
@@
Simulation name central resolution $R_{cv}$ $t_{sf}$
Low res 8 pc 3.0 dx $t_{ff}/0.2$
2x cvr 8 pc 6.0 dx $t_{ff}/0.2$
2x tsf 8 pc 3.0 dx $t_{ff}/0.1$
@@
The development of the outflow appears to be more sensitive to changes in spatial resolution and control volume size than the SFH. The HI outflow rate in the low-resolution run exceeds that of the fiducial run by a factor of 1.5 at z = 500 pc and by a factor of 4 at the edge of the simulation domain. Several factors may contribute to this discrepancy. First, lower resolution leads to SNe energy being injected into a larger volume, potentially resulting in greater radiative cooling losses during the free expansion phase but also more efficient acceleration of surrounding cool gas through direct energy injection. Second, lower resolution may limit the formation of unstable structures, preventing hot, ionized gas from escaping through gaps and channels between clumps, thus enhancing the efficiency of SN-driven outflows. Third, as shown in Section 4, the evolution of outflow rates is not synchronous in different simulations.

### v4

@@ Acknowledgements
We are grateful to the anonymous reviewer for his/her insightful comments and suggestions. We thank the helpful discussions with Antonios, Katsianis and Cheng, Li. This work is supported by the National SKA Program of China (Grant Nos. 2022SKA0110200 and 2022SKA0110202), by the National Natural Science Foundation of China (NFSC) through grants 11733010 and 12173102, and by the China Manned Space Program through its Space Application System. The authors would like to thank National Supercomputer Center in Guangzhou for providing high performance computational resources. Post simulation analysis was performed on the HPC facility of the School of Physics and Astronomy, Sun Yat-Sen University.
@@ A.1 Stellar component
Observations in the 1970s and 1980s had identified M82 as a dusty disk galaxy with a central core. Later observations further revealed the existence of more complex structures. For instance, Telesco et al. (1991) identified characteristics of a bar in the central region of M82. The same structure was confirmed by Wills et al. (2000). In addition, some studies revealed the existence of spiral arms (Mayya et al. 2005) in M82. However, in this work, we omit the effect of bar and spiral arms, only account for a stellar fraction from a bulge, or in some contexts stellar spheroid (Tomisaka and Bregman 1993; Strickland and Stevens 2000), and an exponential disk. The scale height and scale length of both the bulge and stellar disk have been constrained by several observations (Ichikawa et al. 1995; Schreiber et al. 2001;Mayya et al. 2006).
@@
To describe the mass distribution and the potential of the stellar disk component, we use the widely adopted Miyamoto and Nagai (1975) disk profile. The density profile of a Miyamoto-Nagai disk is described by
@@
The bulge overlaps spatially with the recent starburst region. Therefore, a significant portion of the bulge’s stellar mass likely originated from the recent starburst events we aim to simulate. Hence, unlike previous studies, we take a different approach in our simulation by incorporating the bulge as a part of the initial gas disk. By directly adding the mass of the bulge to the initial mass of the gas disk, we can obtain a gas disk profile that more accurately reflects the initial condition before the onset of the recent starburst. This method also allows for a more self-consistent formation of stars in the nuclear region during our starburst simulation. The detailed setup of the gas disk is described in the following section.
Note that there are some uncertainties associated with the chosen parameters for both the bulge and stellar disk. These uncertainties arise from the notable extinction and assumptions on the stellar population, initial mass function, star formation history, etc. A more accurate understanding of the stellar component may be of considerable interest, but is beyond our ability.
@@ A.2 Gas disk
Most of the previous simulations of galactic wind in M82-like galaxies uses a single Miyamoto and Nagai (1975) gas disk (e.g., Strickland and Stevens 2000; Cooper et al. 2008). More recently, Martini et al. (2018) proposed a different setup of the mass model to obtain the analytical form of the gravity potential. In their work, a triple Miyamoto-Nagai disk model (Smith et al. 2015) was employed. Though the Smith et al. (2015) model may produce an accurate fit to the mass distribution of an exponential gas disk, it fails to reproduce the gas profile of M82. Leroy et al. (2015) found the density profile of HI in M82 best fitted by a projected $\rho\propto r^{-2}$ density profile, and $\rm{H_{2}}$ best fitted by a projected $\rho\propto r^{-3.5}$ density profile along the major axis. We employ a similar method to that described by Smith et al. (2015), using three Miyamoto-Nagai disks to fit the projected density profile found by Leroy et al. (2015). Our choice of parameters for the triple Miyamoto-Nagai disk model is given in Table 3.
@@ A.3 Dark matter halo
Early research of the M82’s rotational curve indicated that its rotational curve profile was likely to be Keplerian (Goetz et al. (1990)). Consequently, many previous simulations of galactic wind in M82-like galaxies did not include the gravitational contribution from the dark matter halo. However, more recent observations of the CO line and star clusters suggest a different scenario. Greco et al. (2012) found the rotational curve of M82 is considerably flattened towards the outer region, which is commonly interpreted as evidence for the existence of the dark matter halo (DM halo hereafter). However, the complexity of interactions and mergers among the three members of the M81 group complicates efforts to accurately constrain the parameters of the DM halo. In Oehm et al. (2017), three-body simulation and Markov chain Monte Carlo (MCMC) method are employed to constrain the initial parameters of the three DM halos, and a high-resolution N-body simulation was conducted to further investigate the evolution of the triplet. They demonstrated that the current DM halos, in which M82 and its neighboring galaxies are embedded, arise from the distortion, blending, and merging of the original DM halos. Simple analytical models centered at M82 may struggle to reproduce the true profile of the galaxy’s DM halo. For the sake of simplicity, we adopt a modified NFW profile in this work. The density distribution of the original NFW dark matter halo is given by
@@
Though the NFW profile has been widely adopted, it is not without its shortcomings, particularly in the scenario we are considering. The density of an NFW DM halo does not converge to a finite value towards the center, while observations suggest that many galaxies have a constant-density core in the central area of their DM halos. Such a discrepancy is known as the cusp-core problem. Read et al. (2016) in their simulation shows that star formation and stellar feedback activities occurring in the embedded dwarf galaxy can transform the initial NFW DM halo into a ’core-NFW’ profile with a constant density core in the starburst region ($R_{1/2}$) within a Hubble time. The profile of a ’Core-NFW’ halo can be described by
@@
With the $c-\rm{M_{vir}}$ relation enforced, two free parameters remain: $\rm{M_{vir}}$ and $R_{vir}$. We used the Markov Chain Monte Carlo(MCMC) method to find the optimal combination of the two parameters that best fit the observed rotation curve of Greco et al. (2012). To obtain a more accurate result, we excluded the central 200 pc region, where the velocity dispersion is significantly higher than the local rotational velocity. Additionally, we also excluded the velocity bump located $\sim 500$ pc at the eastern part of M82, which is caused by an SSC designated as cluster ’z’ by McCrady and Graham (2007). We find the best fitting halo mass to be $\rm{M_{vir}}=6.0\times 10^{10}\,\rm{M_{\odot}}$, and the best fitting virial radius is $R_{vir}=53\,\rm{kpc}$. Fig 1 shows that the overall rotation curve of our model closely matches the observed result of Greco et al. (2012).
Appendix B Refinement scheme
Figure 20: The refinement map of our simulations. The resolution is given by $\rm{\Delta x}=64/2^{refinement\ level}$.
In our simulations, we adopted a static mesh refinement scheme. Resolving the evolution of the gas disk, the starburst, and supernova injection ideally requires parsec-scale resolution and even better. Constrained by computational resources, we use a maximum resolution of $\Delta x=4$ pc in the central, high-density region of the gas disk. The resolution then decreases gradually with distance from the center, reaching $64$ pc in the outer regions of the domain. There are four levels of refinement. The refinement map is shown in Figure 20.
Appendix C Convergence test
Figure 21: From the top to the bottom: star formation history, the integrated SFH, and the cumulative cluster mass function in simulation simulation FD, Low res, $\rm{2\times}$ cvr, and $\rm{2\times}$ tsf (see Table 4 for more details).
@@ Appendix C Convergence test
Simulation name central resolution $R_{cv}$ $t_{\rm{sf}}$
Low res 8 pc 1.5 dx $t_{ff}/0.2$
2x cvr 8 pc 3.0 dx $t_{ff}/0.2$
2x tsf 8 pc 1.5 dx $t_{ff}/0.1$
@@
The development of the outflow appears to be more sensitive to changes in spatial resolution and control volume size than the SFH. The HI outflow rate in the low-resolution run exceeds that of the fiducial run by a factor of 1.5 at z = 500 pc and by a factor of 4 at the edge of the simulation domain. Several factors may contribute to this discrepancy. First, lower resolution leads to SNe energy being injected into a larger volume, potentially resulting in greater radiative cooling losses during the free expansion phase but also more efficient acceleration of surrounding cool gas through direct energy injection. Second, lower resolution may limit the formation of unstable structures, preventing hot, ionized gas from escaping through gaps and channels between clumps, thus enhancing the efficiency of SN-driven outflows. Third, as shown in Section 4, the evolution of outflow rates is not synchronized in different simulations.